A veterinarian wishes to use 132 feet of chain-link fencing to enclose a rectangular region and subdivide the region into smaller rectangular regions. If the total enclosed area is 576 square feet, find the dimensions of the enclosed region.
Answer by jorel555(1290)

Let w and l be the width and the length. Then:
2w+2l=132
2l=132-2w
l=66-w
w(66-w)=576
66w-w^2-576=0
w^2-66w+576=0
Using the quadratic formula, we get a length and width of 55.65 ft. X 10.35 ft.!!!!!!!!!!!